is (2,2) a solution of y = 6x -4.
Answer by jim_thompson5910(35256)

y = 6x -4

2 = 6(2)-4 ... Plug in x = 2 and y = 2

2 = 12 - 4

2 = 8 ... which is FALSE

So (2,2) is NOT a solution to y = 6x -4